# Zero-Shot Learning of Causal Models

### v2

@@ E.1 Ablation Study of Encoder
Table 14: Encoder Ablation for Noise Prediction. We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations against two variants. One variant corresponds to the encoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.
@@
Table 15: Encoder Ablation for Sample Generation. We compare Cond-FiP against the baselines for the task of generating samples from input noise variables against two variants. One variant corresponds to the encoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.
@@
Table 16: Encoder Ablation for Interventional Generation. We compare Cond-FiP against the baselines for the task of generating interventional data from input noise variables against two variants. One variant corresponds to the encoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.
E.2 Ablation Study of Decoder
@@
Note that in the main results (Tabel 2Table 3) we show the performances of Cond-FiP trained by sampling SCMs with both linear and non-linear functional relationships. Hence, this ablations helps us to understand whether the strategy of training on mixed functional relationships can bring more generalization to the amortization process, or if we should have trained decoders specialized for linear and non-linear functional relationships.
@@
Table 17: Decoder Ablation for Sample Generation. We compare Cond-FiP for the task of generating samples from input noise variables against two variants. One variant corresponds to a decoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.
@@
Table 18: Decoder Ablation for Interventional Generation. We compare Cond-FiP against two variants for the task of interventional data from input noise variables. One variant corresponds to a decoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario.
E.3 Sensitivity to Distribution Shifts
In our main results (Table 123) we experimented with datasets sampled from SCM following a different distribution (LIN OUT, RFF OUT) than the datasets used for training Cond-FiP (LIN IN, RFF IN). As expected, the performance of all methods drop in the presence of distribution shift. We now analyze how sensitive is Cond-FiP to distribution shifts by comparing its performance across scenarios as the severity of the distribution shift is increased.
To illustrate how we control the magnitude of distribution shift, we discuss the difference in the distribution of causal mechanisms across ${\mathbb{P}}_{\text{IN}}$ and ${\mathbb{P}}_{\text{OUT}}$. The distribution shift arises because the support of the parameters of causal mechanisms changes from ${\mathbb{P}}_{\text{IN}}$ to ${\mathbb{P}}_{\text{OUT}}$, as detailed below.
  • Linear Functional Mechanism.
    • In-Distribution (${\mathbb{P}}_{\text{IN}}$)
        Weights: $\sim U_{\pm}(1,3)$, Bias $\sim U(-3,3)$.
    • Out-of-Distribution (${\mathbb{P}}_{\text{OUT}}$)
        Weights: $\sim U_{\pm}(0.5,2)\cup U_{\pm}(2,4)$, Bias $\sim U(-3,3)$.
  • RFF Functional Mechanism.
    • In-Distribution (${\mathbb{P}}_{\text{IN}}$)
    • Out-of-Distribution (${\mathbb{P}}_{\text{OUT}}$):
For example, for linear causal mechanism case, the weights in ${\mathbb{P}}_{\text{IN}}$ are sampled uniformly from $(-3,-1)\cup(1,3)$; while in ${\mathbb{P}}_{\text{OUT}}$ they are sampled from uniformly from $(0.5,4)$. We now change the support set of the parameters in ${\mathbb{P}}_{\text{OUT}}$ to $(0.5\alpha,4\alpha)$, so that by increasing $\alpha$ we make the distribution shift more severe. We follow this procedure for the support set of all the parameters associated with functional mechanisms and generate distributions (${\mathbb{P}}_{\text{OUT}}(\alpha)$) with varying shift w.r.t ${\mathbb{P}}_{\text{IN}}$ by changing $\alpha$. Note that $\alpha=1$ corresponds to the same ${\mathbb{P}}_{\text{OUT}}$ as the one used for sampling datasets in our main results.
Tables 1920, and 21 provide results for the case of controlling shift via causal mechanisms, for the task of noise prediction, sample generation, and interventional generation respectively. We find that the performance of Cond-FiP does not change much as we increase $\alpha$, indicating that Cond-FiP is robust to the varying levels of distribution shits in causal mechanisms.
However, for the case of controlling shift via noise variables (Table 2223, and 24) we find that Cond-FiP is quite sensitive to the varying levels of distribution shift in noise variables. The performance of Cond-FiP degrades with increasing magnitude of the shift ($\alpha$) for all the tasks.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.07$ $(0.01)$ $0.10$ $(0.01)$
10 2 $0.06$ $(0.01)$ $0.10$ $(0.01)$
10 5 $0.05$ $(0.01)$ $0.10$ $(0.01)$
10 10 $0.05$ $(0.01)$ $0.10$ $(0.01)$
20 1 $0.07$ $(0.0)$ $0.12$ $(0.0)$
20 2 $0.06$ $(0.0)$ $0.13$ $(0.01)$
20 5 $0.05$ $(0.0)$ $0.11$ $(0.01)$
20 10 $0.05$ $(0.0)$ $0.10$ $(0.01)$
50 1 $0.07$ $(0.01)$ $0.14$ $(0.01)$
50 2 $0.05$ $(0.01)$ $0.17$ $(0.01)$
50 5 $0.05$ $(0.01)$ $0.14$ $(0.01)$
50 10 $0.04$ $(0.0)$ $0.14$ $(0.01)$
100 1 $0.07$ $(0.01)$ $0.16$ $(0.01)$
100 2 $0.05$ $(0.01)$ $0.18$ $(0.0)$
100 5 $0.05$ $(0.0)$ $0.17$ $(0.01)$
100 10 $0.05$ $(0.0)$ $0.16$ $(0.01)$
Table 19: Results for Noise Prediction under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the case in main results Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.05$ $(0.01)$ $0.08$ $(0.01)$
10 2 $0.05$ $(0.0)$ $0.07$ $(0.01)$
10 5 $0.05$ $(0.0)$ $0.07$ $(0.01)$
10 10 $0.06$ $(0.0)$ $0.06$ $(0.01)$
20 1 $0.07$ $(0.01)$ $0.30$ $(0.03)$
20 2 $0.06$ $(0.01)$ $0.34$ $(0.05)$
20 5 $0.06$ $(0.01)$ $0.35$ $(0.05)$
20 10 $0.06$ $(0.01)$ $0.29$ $(0.07)$
50 1 $0.07$ $(0.0)$ $0.48$ $(0.07)$
50 2 $0.07$ $(0.0)$ $0.47$ $(0.07)$
50 5 $0.07$ $(0.01)$ $0.38$ $(0.06)$
50 10 $0.07$ $(0.01)$ $0.32$ $(0.06)$
100 1 $0.09$ $(0.01)$ $0.57$ $(0.07)$
100 2 $0.09$ $(0.01)$ $0.60$ $(0.05)$
100 5 $0.09$ $(0.01)$ $0.58$ $(0.05)$
100 10 $0.12$ $(0.02)$ $0.56$ $(0.06)$
Table 20: Results for Sample Generation under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the case in main results Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.07$ $(0.01)$ $0.11$ $(0.01)$
10 2 $0.07$ $(0.01)$ $0.11$ $(0.01)$
10 5 $0.07$ $(0.01)$ $0.10$ $(0.01)$
10 10 $0.06$ $(0.01)$ $0.10$ $(0.01)$
20 1 $0.14$ $(0.03)$ $0.31$ $(0.03)$
20 2 $0.10$ $(0.02)$ $0.33$ $(0.04)$
20 5 $0.17$ $(0.1)$ $0.34$ $(0.04)$
20 10 $0.10$ $(0.03)$ $0.28$ $(0.05)$
50 1 $0.12$ $(0.02)$ $0.48$ $(0.07)$
50 2 $0.12$ $(0.03)$ $0.47$ $(0.07)$
50 5 $0.11$ $(0.01)$ $0.39$ $(0.06)$
50 10 $0.11$ $(0.02)$ $0.32$ $(0.06)$
100 1 $0.14$ $(0.02)$ $0.58$ $(0.07)$
100 2 $0.13$ $(0.02)$ $0.60$ $(0.06)$
100 5 $0.14$ $(0.03)$ $0.58$ $(0.05)$
100 10 $0.18$ $(0.04)$ $0.55$ $(0.06)$
Table 21: Results for Interventional Generation under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the case in main results Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.07$ $(0.01)$ $0.10$ $(0.01)$
10 2 $0.07$ $(0.01)$ $0.11$ $(0.01)$
10 5 $0.07$ $(0.01)$ $0.18$ $(0.02)$
10 10 $0.08$ $(0.01)$ $0.26$ $(0.04)$
20 1 $0.07$ $(0.0)$ $0.12$ $(0.0)$
20 2 $0.07$ $(0.0)$ $0.16$ $(0.01)$
20 5 $0.07$ $(0.0)$ $0.30$ $(0.01)$
20 10 $0.07$ $(0.0)$ $0.41$ $(0.02)$
50 1 $0.07$ $(0.01)$ $0.14$ $(0.01)$
50 2 $0.07$ $(0.01)$ $0.19$ $(0.01)$
50 5 $0.07$ $(0.01)$ $0.33$ $(0.02)$
50 10 $0.07$ $(0.01)$ $0.44$ $(0.02)$
100 1 $0.07$ $(0.01)$ $0.16$ $(0.01)$
100 2 $0.07$ $(0.01)$ $0.22$ $(0.0)$
100 5 $0.07$ $(0.01)$ $0.35$ $(0.01)$
100 10 $0.07$ $(0.01)$ $0.44$ $(0.01)$
Table 22: Results for Noise Prediction under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the case in main results Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.05$ $(0.01)$ $0.08$ $(0.01)$
10 2 $0.05$ $(0.0)$ $0.13$ $(0.03)$
10 5 $0.05$ $(0.01)$ $0.28$ $(0.06)$
10 10 $0.05$ $(0.01)$ $0.36$ $(0.08)$
20 1 $0.07$ $(0.01)$ $0.30$ $(0.03)$
20 2 $0.07$ $(0.01)$ $0.45$ $(0.04)$
20 5 $0.07$ $(0.01)$ $0.59$ $(0.03)$
20 10 $0.07$ $(0.01)$ $0.58$ $(0.02)$
50 1 $0.07$ $(0.0)$ $0.48$ $(0.07)$
50 2 $0.07$ $(0.0)$ $0.59$ $(0.06)$
50 5 $0.07$ $(0.0)$ $0.64$ $(0.03)$
50 10 $0.07$ $(0.0)$ $0.58$ $(0.02)$
100 1 $0.09$ $(0.01)$ $0.57$ $(0.07)$
100 2 $0.09$ $(0.01)$ $0.63$ $(0.05)$
100 5 $0.09$ $(0.01)$ $0.65$ $(0.03)$
100 10 $0.09$ $(0.01)$ $0.59$ $(0.02)$
Table 23: Results for Sample Generation under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the case in main results Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.07$ $(0.01)$ $0.11$ $(0.01)$
10 2 $0.07$ $(0.01)$ $0.14$ $(0.02)$
10 5 $0.07$ $(0.01)$ $0.25$ $(0.05)$
10 10 $0.07$ $(0.01)$ $0.32$ $(0.06)$
20 1 $0.14$ $(0.03)$ $0.31$ $(0.03)$
20 2 $0.14$ $(0.03)$ $0.42$ $(0.03)$
20 5 $0.14$ $(0.03)$ $0.57$ $(0.03)$
20 10 $0.14$ $(0.03)$ $0.56$ $(0.02)$
50 1 $0.12$ $(0.02)$ $0.48$ $(0.07)$
50 2 $0.12$ $(0.01)$ $0.58$ $(0.06)$
50 5 $0.12$ $(0.01)$ $0.65$ $(0.04)$
50 10 $0.12$ $(0.01)$ $0.59$ $(0.02)$
100 1 $0.14$ $(0.02)$ $0.58$ $(0.07)$
100 2 $0.14$ $(0.02)$ $0.65$ $(0.06)$
100 5 $0.14$ $(0.02)$ $0.67$ $(0.04)$
100 10 $0.14$ $(0.02)$ $0.60$ $(0.03)$
Table 24: Results for Interventional Generation under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the case in main results Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
Appendix F Comparing Cond-FiP with CausalNF
We also compare Cond-FiP with CausalNF (Javaloy et al., 2023) for the task of noise prediction (Table 25) and sample generation (Table 26). The test datasets consist of $n_{\text{test}}=400$ samples, exact same setup as in our main results (Table 1, Table 2, and Table 3). To ensure a fair comparison, we provided CausalNF with the true causal graph.

### v3

@@ G.1 Ablation Study of Encoder
Table 19: Encoder Ablation for Noise Prediction. We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations against two variants. One variant corresponds to the encoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Results show that training on only non-linear SCMs (Cond-FiP(RFF)) gives similar performance as training on both linear and non-linear SCMs (Cond-FiP).
@@
Table 20: Encoder Ablation for Sample Generation. We compare Cond-FiP against the baselines for the task of generating samples from input noise variables against two variants. One variant corresponds to the encoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Results show that training on only non-linear SCMs (Cond-FiP(RFF)) gives similar performance as training on both linear and non-linear SCMs (Cond-FiP).
@@
Table 21: Encoder Ablation for Interventional Generation. We compare Cond-FiP against the baselines for the task of generating interventional data from input noise variables against two variants. One variant corresponds to the encoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Results show that training on only non-linear SCMs (Cond-FiP(RFF)) gives similar performance as training on both linear and non-linear SCMs (Cond-FiP).
G.2 Ablation Study of Decoder
@@
Note that in the main results (Table 2Table 3) we show the performances of Cond-FiP trained by sampling SCMs with both linear and non-linear causal mechanisms. Hence, this ablations helps us to understand whether the strategy of training on mixed causal mechanisms can bring more generalization to the amortization process, or if we should have trained decoders specialized for linear and non-linear functional relationships.
@@
Table 22: Decoder Ablation for Sample Generation. We compare Cond-FiP for the task of generating samples from input noise variables against two variants. One variant corresponds to a decoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Results indicate that training on both linear and non-linear SCMs is crucial to generalize effectively in all scenarios.
@@
Table 23: Decoder Ablation for Interventional Generation. We compare Cond-FiP against two variants for the task of interventional data from input noise variables. One variant corresponds to a decoder trained on SCMs with only linear functional relationships, Cond-FiP(LIN). Similarly, we have another variant where the decoder was trained on SCMs with only rff functional relationships, Cond-FiP(RFF). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Results indicate that training on both linear and non-linear SCMs is crucial to generalize effectively in all scenarios.
Appendix H Experiments on CSuite with Complex Noise Distributions
(a) Noise Prediction
(b) Sample Generation
(c) Interventional Generation
Figure 7: We compare Cond-FiP against the baselines for the different evaluation tasks on the Large Backdoor and Weak Arrow datasets from the CSuite benchmark, where the noise distribution is modified to be a multi-modal gaussian mixture model. We experiment with $6$ different cases of the noise distribution for each dataset. The y-axis denotes the RMSE for the respective tasks across the $12$ scenarios (datasets & noise distribution). Results indicate that Cond-FiP can generalize to instances with more complex noise distributions like gaussian mixture models.
To conduct more OOD evaluations, we modify the noise distribution of the Large Backdoor and Weak Arrow datasets from the Csuite benchmark such that the noise variables are sampled from a guassian mixture model (GMM). We considered the following cases for the GMM noise distribution.
  • Noise is sampled with equal probability from either $N(-2,1)$ and $N(2,1)$.
  • Noise is sampled with equal probability from either $N(-2,2)$ and $N(2,2)$.
  • Noise is sampled with equal probability from either $N(-2,1)$ and $N(2,2)$.
  • Noise is sampled with equal probability from either $N(-5,1)$ and $N(5,1)$.
  • Noise is sampled with equal probability from either $N(-5,2)$ and $N(5,2)$.
  • Noise is sampled with equal probability from either $N(-5,1)$ and $N(5,2)$.
This leads to a total of $12$ experimental setting with $6$ different GMM noise distribution for both the Large Backdoor and Weak Arrow datasets from the CSuite benchmark. Results in Figure 7 demonstrate that Cond-FiP remains competitive with baselines across all tasks. Importantly, while baselines were trained from scratch for each specific gaussian mixture noise distribution, Cond-FiP was pretrained only on gaussian noise and generalizes effectively to settings with GMM noise distribtion.
Appendix I Experiments on Real World Benchmark
Method MMD($\widehat{D^{\text{query}}_{\bm{X}}}$, $D^{\text{query}}_{\bm{X}}$) MMD($\widehat{D^{\text{context}}_{\bm{X}}}$, $D^{\text{query}}_{\bm{X}}$) MMD($D^{\text{context}}_{\bm{X}}$, $D^{\text{query}}_{\bm{X}}$)
DoWhy $0.015$ $0.014$ $0.005$
DECI $0.014$ $0.005$ $0.005$
FiP $0.015$ $0.005$ $0.005$
Cond-FiP $0.013$ $0.005$ $0.005$
Table 24: Results for Sachs dataset. We benchmark Cond-FiP against the baselines for the task of generating observational data on the real world Sachs benchmark. Each cell reports the MMD, and we also report the reconstruction error for all of the methods. Results indicate that Cond-FiP matches the performance of baselines trained from scratch.
We use the real world flow cytometry dataset [Sachs et al., 2005] to benchmark Cond-FiP againts the baselines. This dataset contains $n\simeq 800$ observational samples expressed in a $d=11$ dimensional space, and the reference (true) causal graph. We split this into context $D^{\text{context}}_{\bm{X}}\in{\mathbb{R}}^{n_{\text{context}}\times d}$ and queries $D^{\text{query}}_{\bm{X}}\in{\mathbb{R}}^{n_{\text{query}}\times d}$, each of size $n_{\text{context}}=n_{\text{query}}=400$. Note that the context dataset is to used to train the baselines and obtain dataset embedding for Cond-FiP, while the query dataset is used for evaluation of all the methods.
Since we don’t have access to the true causal mechanisms, we cannot compute RMSE for noise prediction or sample generation like we did in our experiments with synthetic benchmarks. Instead for each method, we obtain the noise predictions $\widehat{D^{\text{context}}_{\bm{N}}}$ on the context, and use it to fit a gaussian distribution for each component (node). Then we use the learned gaussian distribution to sample new noise variables, $\widehat{D^{\text{query}}_{\bm{N}}}$, which are mapped to the observations as per the causal mechanisms learned by each method, $\widehat{D^{\text{query}}_{\bm{X}}}$. Finally, we compute the maximum mean discrepancy (MMD) distance between $\widehat{D^{\text{query}}_{\bm{X}}}$ and $D^{\text{query}}_{\bm{X}}$ as metric to determine whether the method has captured the true causal mechanisms. For consistency, we also evaluate the reconstruction performances of the models by using directly the inferred noise from context $\widehat{D^{\text{context}}_{\bm{N}}}$ from the models, and then compute MMD between their reconstructed data ($\widehat{D^{\text{context}}_{\bm{X}}}$) and the query data ($D^{\text{query}}_{\bm{X}}$).
Table 24 presents our results, where for reference we also report the MMD distance between samples from the context and query split, which should serve as the gold standard since both the datasets are sampled from the same distribution. We find that Cond-FiP is competitive with the baselines that were trained from scratch. Except DoWhy, the MMD distance with reconstructed samples from the methods are close to oracle performance.
No Interventional Generation Results. Note that Cond-FiP (and the other baselines considered in this work) only supports hard interventions while the interventional data available for Sachs are soft interventions (i.e. the interventional operations applied are unknown). Hence, we are unable to provide a comprehensive evaluation of Cond-FiP (as well as the other baselines) for interventional predictions on Sachs.
Appendix J Comparing Cond-FiP with CausalNF
We also compare Cond-FiP with CausalNF [Javaloy et al., 2023] for the task of noise prediction (Table 25) and sample generation (Table 26). The test datasets consist of $n_{\text{test}}=400$ samples, exact same setup as in our main results (Table 1, Table 2, and Table 3). To ensure a fair comparison, we provided CausalNF with the true causal graph.